x -1 3 5 8 9

p(x) 0.2 0.1 0.4 0.2 0.1

1 list the values that x may assume (use ascending order)?

2 what value of x is most probable?

3 what is the probability that x is greater than 0?

4. what is the probability that x=1?
Answer #1

solution

1) ascending order x value = -1 3 5 8 9

2) x=3,5

3)P(x > 0)= p(x=3)+p(x=5)+p(x=8)+p(x=9)

=0.1+0.4+0.2+0.1

=0.8

probability=0.8

4)p(x=1) =0.0